OpenStudy (anonymous):

-7/8t<-14 solve then write as interval

OpenStudy (anonymous):

t>16

OpenStudy (saifoo.khan):

(16, infinty)

OpenStudy (anonymous):

t<1/16 \[(-\infty,\frac{1}{16})\]

OpenStudy (saifoo.khan):

it must be: \[-\frac{7}{8} t < -14\]?

OpenStudy (anonymous):

no, larger than 16

OpenStudy (anonymous):

okay thanks i was just checking. I just didn't know how to put them in interval

OpenStudy (anonymous):

Where is the t krissy?

OpenStudy (anonymous):

on the less than side of -7/8